OpenStudy (taylormarie213):

The length of a rectangle is three more than twice the width. If the perimeter is 72 inches, find the dimensions of the rectangle.

OpenStudy (wolf1728):

2L = 3W 2L + 2W = 72 Which solves to L = 21.6 and W = 14.4 Testing the answers 2*21.6 = 3*14.4 2*21.6 + 2*14.4 = 43.2 + 28.8 = 72 The answer doesn't have be integers.

OpenStudy (taylormarie213):

thanks!

OpenStudy (wolf1728):

Wow - I see this has been answered in another location and it has integer answers. Here's the link: http://openstudy.com/study#/updates/522b94aee4b0a8663d55a828